OpenStudy (anonymous):

Write an equation of a line that passes through the points (9,-5) (3,-5) I know the slope equation is y=mx+b with m=slope the slope of this equation is 5/3 so y=5/3x+b but how do I find b?

OpenStudy (anonymous):

you find b by substituting the x and y from either (9,-5) or (3,-5) like y=5/3x+b -5=5/3(9)+b solve.

OpenStudy (anonymous):

oh, okay! thanks so much! :)

OpenStudy (anonymous):

no problem